# On the Impacts of Contexts on Repository-Level Code Generation

### v3

@@ Abstract
CodeLLMs have gained widespread adoption for code generation tasks, yet their capacity to handle repository-level code generation with complex contextual dependencies remains underexplored. Our work underscores the critical importance of leveraging repository-level contexts to generate executable and functionally correct code. We present RepoExec, a novel benchmark designed to evaluate repository-level code generation, with a focus on three key aspects: executability, functional correctness through comprehensive test case generation, and accurate utilization of cross-file contexts. Our study examines a controlled scenario where developers specify essential code dependencies (contexts), challenging models to integrate them effectively. Additionally, we introduce an instruction-tuned dataset that enhances CodeLLMs’ ability to leverage dependencies, along with a new metric, Dependency Invocation Rate (DIR), to quantify context utilization. Experimental results reveal that while pretrained LLMs demonstrate superior performance in terms of correctness, instruction-tuned models excel in context utilization and debugging capabilities. RepoExec offers a comprehensive evaluation framework for assessing code functionality and alignment with developer intent, thereby advancing the development of more reliable CodeLLMs for real-world applications. The dataset and source code are available at https://github.com/FSoft-AI4Code/RepoExec.
@@ 1 Introduction
Code Large Language Models (CodeLLMs) have emerged as powerful tools for assisting with coding tasks Wang et al. (2021, 2023); Feng et al. (2020); Allal et al. (2023); Li et al. (2023); Lozhkov et al. (2024); Guo et al. (2024); Pinnaparaju et al. (2024); Zheng et al. (2024); Roziere et al. (2023); Nijkamp et al. (2022); Luo et al. (2023); Xu et al. (2022); Bui et al. (2023). While these models excel at generating code from natural language requirements or completing individual lines of code, their application in real-world, professional software development scenarios presents more complex challenges. A critical aspect of this complexity lies in leveraging relevant contexts across an entire software repository, which raises two pivotal questions. First, to what extent are the retrieved dependencies accurate and relevant, rather than potential noise in the input? Second, how effectively do LLMs process and incorporate the provided dependencies into their generated code? These inquiries are central to understanding the capabilities and limitations of CodeLLMs in repository-level code generation, where completing a single line of code might require making API calls to functions within the same file (in-file context) or across different files (cross-file context).
Existing repository-level code generation benchmarks, such as RepoBench Liu et al. (2023b), RepoCoder Zhang et al. (2023a), CrossCodeEval Ding et al. (2023), CoCoMIC Ding et al. (2024), and DevEval Li et al. (2024), have made significant strides in assessing code generation at a repository-level scale. However, these benchmarks face several limitations that hinder their ability to comprehensively evaluate real-world coding scenarios:
@@
    Lack of an executable environment: This leads to reliance on match-based metrics, which are not robust for assessing the functional correctness of generated code.
  2. 2.
    Inadequate control over test quality: Although RepoCoder and DevEval ensure executability, they depend on pre-existing test cases extracted from repositories, limiting the robustness of evaluation and hindering the scalability of the data pipeline Liu et al. (2023a).
  3. 3.
    Overreliance on functional correctness metrics like pass@k: These metrics, inherited from standalone code generation studies Chen et al. (2021); Austin et al. (2021), are insufficient in the context of repository-level code generation.
Moreover, real-world development practices often involve breaking down large functions into smaller modules (functions, classes, or variables) to enhance problem-solving efficiency. Consequently, generated code must effectively call these predefined dependencies to align with developer intent. However, LLMs may produce code that passes test cases but relies on inefficient workarounds or unnecessarily reimplements predefined functions. This can lead to suboptimal implementations, code duplication, and potential risks such as technical debt and code smell issues (Maldonado and Shihab, 2015; Sierra et al., 2019; Rasool and Arshad, 2015; Santos et al., 2018).
Given these challenges, it is evident that evaluating CodeLLMs based solely on functional correctness is insufficient. There is a pressing need for a more comprehensive evaluation strategy that assesses not only the correctness of generated code but also its alignment with developer intent, efficiency in utilizing provided dependencies, and adherence to best practices in software development. Crucially, our research reveals a strong correlation between functional correctness (as measured by pass@1) and effective dependency utilization, suggesting that models that better leverage contextual information tend to produce more accurate and functionally correct code.
To effectively address these shortcomings, we introduce a paradigm for evaluation that emphasizes both functional correctness and dependency utilization, while also exploring their relation. Specifically, we introduce the Dependency Invocation Rate (DIR), a novel metric that measures the proportion of code dependencies successfully integrated into the generated code. Higher DIR values indicate better model comprehension of dependencies, while lower values suggest challenges in dependency utilization. Importantly, our findings demonstrate a strong correlation between DIR and pass@1 rates, underscoring the critical role of effective context utilization in producing high-quality, functionally correct code.
We also introduce RepoExec, a novel benchmark specifically designed to overcome the limitations of existing benchmarks, providing a robust and comprehensive evaluation for code generation at real-world applicability. In detail, RepoExec advances with two primary focuses:
  • Executability and Dynamic Test Case Generation: RepoExec ensures an executable environment within the repository context, thereby providing a reliable evaluation of functional correctness. Besides, it also incorporates a practical mechanism for dynamically generating high-coverage test cases tailored to the functionality of the newly generated code, ensuring that the code performs its intended tasks accurately.
  • Dependency Usage: RepoExec provides a pipeline to evaluate the usage of dependencies across CodeLLMs, offering insights on how well code dependencies should be provided. Additionally, we introduce an instruction-tuning dataset with dependency calls, which significantly improves CodeLLMs’ ability to leverage code dependencies and produce better results.
Our experiments with RepoExec reveal key insights into repository-level code generation by CodeLLMs. Models achieve optimal performance with full context dependencies, though smaller contexts outperform medium ones due to formatting issues in the BasePrompt. While pretrained LLMs like Codellama-34b-Python lead in pass@1 rates, instruction-tuned LLMs are better at managing dependencies, despite sometimes generating overly complex code. In contrast, pretrained LLMs produce accurate code but struggle with dependency utilization. These findings further reinforce the observed correlation between pass@1 and DIR, highlighting the importance of balancing accuracy and context utilization in CodeLLM performance.
Furthermore, we explore two potential approaches for enhancing the quality of generated code: Multi-round debugging and Instruction tuning. Multi-round debugging with test execution, especially with GPT-3.5 and WizardCoder, significantly improves both pass@1 scores and DIR after three rounds. Fine-tuning with our dependency-enhanced dataset not only boosts pass@1 and DIR metrics but also minimizes computational costs, highlighting the importance of executable code testing and the advantages of instruction-tuning and multi-round debugging in enhancing CodeLLMs’ performance and dependency management.
In summary, our contributions are as follows:
  1. 1.
    We introduce a novel evaluation paradigm aimed at providing a robust and comprehensive assessment for repository-level code generation. Our approach not only evaluates the functional correctness of generated code but also considers additional quality factors such as maintainability and adherence to clean code principles through effective dependency utilization.
@@
    We introduce RepoExec, a novel benchmark that aligns with our evaluation paradigm, addressing the gaps in existing benchmarks and offering a comprehensive assessment of code generation quality.
@@
    We implement an effective pipeline within RepoExec from dependency extraction, dynamically generate high-coverage test cases and automatic evaluation with execution and code dependencies. Our pipeline offers practical usage and scalability for the community.
@@
    We release a tool named pydepcall [^1] https://github.com/FSoft-AI4Code/pydep/tree/main to extract dependencies for all functions within any repository, providing a practical use for advancing research in this domain.
  6. 6.
    Our experiments provide crucial insights into the performance of CodeLLMs in repository-level code generation. We demonstrate that while foundation models achieve high initial accuracy, instruction-tuned models excel in dependency management and addressing complex scenarios. Additionally, our multi-round debugging tests show notable improvements in model performance, particularly with enhancements in dependency management. Importantly, we establish a strong correlation between pass@1 and DIR, emphasizing that models which better utilize contexts tend to produce more functionally correct code.
@@ 2 Related works
Coding-related tasks have been crucial for assessing the performance of Large Language Models (LLMs), with code generation emerging as a primary focus (Chen et al., 2021; Li et al., 2023; Jiang et al., 2024; Touvron et al., 2023; Roziere et al., 2023; Xu et al., 2022; Allal et al., 2023; Nijkamp et al., 2022). Early benchmarks have been introduced to address this issue Yin et al. (2018); Iyer et al. (2018); Nguyen et al. (2023); Chen et al. (2021); Austin et al. (2021); Hendrycks et al. (2021); however, they often had limited scope or employed weak evaluation approaches. Some benchmarks Yin et al. (2018); Iyer et al. (2018); Nguyen et al. (2023) exhibit domain diversity akin to real-world applications; however, their evaluation methodologies are constrained to match-based metrics, thereby decreasing the reliability of these benchmarks Chen et al. (2021). Meanwhile, benchmarks with reliable evaluation approaches such as HumanEval Chen et al. (2021), MBPP Austin et al. (2021) and APPS Hendrycks et al. (2021) often entail limitations to specific domains like competitive programming. Recently, there have been efforts to extend the domains of generation tasks in various benchmarks. For instance, ExeDS Huang et al. (2022) focuses on data science code generation, while ODEX Wang et al. (2022) serves as an open-domain benchmark for code generation. Besides, all the benchmarks mentioned primarily focus on standalone function generation, lacking consideration for cross-contextual and dependency invocation scenarios.
@@ 3 Evaluation Paradigm
In this section, we outline our approach to achieve a more robust and comprehensive evaluation of repository-level code generation. Our paradigm encompasses two key attributes: Functional correctness and Dependency utilization.
Functional correctness: This evaluation criterion ensures that the code accurately performs its intended tasks and requirements. Specifically, it typically involves using test cases to compare the execution output of the generated code against the expected output for a given input. This criterion has been widely employed for evaluating code generation in numerous studies (Chen et al., 2021; Austin et al., 2021; Zhuo et al., 2024; Li et al., 2024). We follow the well-known automatic metric Pass@k Chen et al. (2021) to measure the functional correctness of generation outputs.
Dependency utilization: However, functional correctness alone cannot fully assess code quality, as it may overlook issues like poor implementation, workaround solutions, or code redundancy. Assuming that human-written solutions are optimal, match-based metrics like BLEU and edit similarity can assess the alignment between the generated code and the high-quality reference solution at the token, word, or character level. However, not all words and tokens contribute to the quality of the code and alternative implementations can preserve the quality despite low similarity scores. For instance, local variable names or comments can vary widely without impacting code quality. Additionally, statements like while and for can be interchangeable while preserving the algorithm’s complexity. Thus, these match-based metrics are also not reliable for accurately measuring code quality. Meanwhile, tokens representing called dependencies—such as third-party packages, functions, variables, and classes within the repository—demonstrate effective use of quality code for efficient implementation. In contrast, ignoring these dependencies may suggest workaround implementations misaligned with human intent, leading to extensive verification and maintenance costs.
Therefore, to assess the models’ ability to utilize provided dependencies in accordance with human intent, we introduce the Dependency Invocation Rate (DIR). This metric represents the percentage of invoked dependencies out of the total number of dependencies provided. Let $D_{g}$ denote the set of identifiers in the generated output, and $D_{s}$ denote the set of provided dependencies extracted from the solution. The Dependency Invocation Rate (DIR) is calculated as follows:
@@
A higher DIR indicates that the model successfully incorporates a larger proportion of the provided dependencies into the generated code, demonstrating a better understanding of the dependencies’ relevance and their intended usage. Conversely, a lower DIR suggests that the model may struggle to identify and utilize the appropriate dependencies, potentially generating code that is less aligned with the human-written solution.
In summary, we state that achieving a high-quality solution requires the generated code to excel in both functional correctness and dependency utilization. Otherwise, it indicates a lack of ability to generate correct solutions for the intended tasks or suggests poor implementation practices, including technical debt or code smell issues.
@@ 4.1 Data Source
Developing an executable benchmark within repository contexts is challenging due to complex setup requirements and frequent lack of clear installation guidelines in repositories. Previous studies (Ding et al., 2024; Shrivastava et al., 2023; Ding et al., 2023; Liao et al., 2023) have often relied on match-based metrics for evaluation, which may not fully capture code functionality. In addition, test cases are essential for assessing code functionality. However, extracting test cases from repositories (Zhang et al., 2023b; Li et al., 2024; Zhang et al., 2024) often relies on available functions and heuristic rules, limiting adaptability and excluding data without existing tests. For instance, Li et al. 2024 found that over 99% of functions were discarded due to the absence of suitable tests. We propose a dataset collection pipeline to ensure that repositories can build executable environments and that test cases are automatically generated.
@@ 4.2 Functions and Dependencies Extraction
We employ static analysis to identify dependencies, excluding identifiers that are function parameters or typing objects. Each dependency name is then mapped to its definition within the repository using a repository graph and static analysis tools. We release our tool pydepcall for community usage and provide a brief description in Appendix H.
@@ 4.3 Test case generation
To overcome the limitation of requiring available tests in the repository for evaluation (as mentioned in Section 4.1), we leverage large language models (LLMs) to generate test cases automatically. Our proposed approach ensures the correctness of created test cases while also controlling and enhancing test coverage. To execute and validate the generated test cases, it is necessary to configure each repository to create an executable environment. Follow Lemieux et al. 2023, we use pipreqs [^2] https://github.com/bndr/pipreqs to identify each project’s dependencies.
In our test generation process, we conduct two phases corresponding to Correctness Control and Coverage Enhancement. After the test generation process, we exclude samples with a line coverage lower than 40%, as they are of insufficient quality to accurately assess the correctness of the generated code.
@@ 4.3.1 Correctness Control
In this stage, we present the procedure for generating initial test cases and ensure the tests’ correctness. Specifically, we use CodeLlama-13b Roziere et al. (2023) and provide the model with the prompt detailed in Appendix C. After generation, we extract the first 20 assertion statements to construct the test cases. To ensure their correctness, we employ syntax and execution-based filtering methods to exclude low-quality test cases.
@@
We use pytest [^3] https://github.com/pytest-dev/pytest to execute the generated tests. Tests are discarded if their execution output displays an error, except for AssertionError. If an assertion error is observed, we try to fix it using the Assertion Fixer.
  • Assertion Fixer: The fixer resolves this error by executing the test case with the target function, extracting the output, and reassigning it to the assertion statement. Given that the return type can be complex in the repository (e.g., a defined object), we use pickle [^4] https://docs.python.org/3/library/pickle.html to preserve the execution results.
  • Multiple time execution: We witness some flaky tests, which produce inconsistent outcomes upon multiple executions due to the inherent randomness in their implementation. To eliminate these instances, we execute each test 10 times to compare the outputs.

### v4

@@ Abstract
CodeLLMs are widely used for code generation, yet their ability to handle repository-level dependencies remains underexplored. We introduce RepoExec, a benchmark for evaluating repository-level code generation, focusing on executability, functional correctness, and dependency utilization. Our study evaluates 18 models, revealing that retaining full dependency context yields the best performance, while smaller context sizes can be misleading. Pretrained LLMs excel in correctness but often reimplement dependencies, while instruction-tuned models better utilize dependencies but sometimes introduce unnecessary complexity. We propose an instruction-tuning dataset that improves dependency handling and introduce a new metric, Dependency Invocation Rate (DIR), to measure context utilization. Experiments show that instruction-tuned models improve DIR by over 10%, and multi-round debugging further enhances both correctness and dependency use. RepoExec provides a comprehensive framework to advance CodeLLMs for real-world applications. The dataset and source code are available at https://github.com/FSoft-AI4Code/RepoExec.
@@ 1 Introduction
Code Large Language Models (CodeLLMs) have emerged as powerful tools for assisting with coding tasks Wang et al. (2021, 2023); Feng et al. (2020); Allal et al. (2023); Li et al. (2023); Lozhkov et al. (2024); Guo et al. (2024); Pinnaparaju et al. (2024); Zheng et al. (2024); Roziere et al. (2023); Nijkamp et al. (2022); Luo et al. (2023); Xu et al. (2022); Bui et al. (2023); Manh et al. (2024); Liu et al. (2024a). While these models excel at generating code from natural language requirements or completing individual lines of code, their application in real-world, professional software development scenarios presents more complex challenges. A critical aspect of this complexity lies in leveraging relevant contexts/dependencies (e.g., function calls, imports, class hierarchy) across an entire software repository, which raises two pivotal questions. First, to what extent are the retrieved contexts accurate and relevant, rather than potential noise in the input? Second, how effectively do LLMs process and incorporate the provided dependencies into their generated code? These inquiries are central to understanding the capabilities and limitations of CodeLLMs in repository-level code generation, where completing a single line of code might require making API calls to functions within the same file (in-file context) or across different files (cross-file context).
Existing repository-level code generation benchmarks, such as RepoBench Liu et al. (2023b), RepoCoder Zhang et al. (2023a), CrossCodeEval Ding et al. (2023), CoCoMIC Ding et al. (2024), and DevEval Li et al. (2024), have advanced the evaluation of code generation at a repository level. However, they exhibit key limitations: (1) Lack of an executable environment, leading to dependence on match-based metrics that fail to assess functional correctness adequately; (2) Insufficient control over unit test quality, reducing the data pipeline scalability and robustness of the evaluation; and (3) Overemphasis on functional correctness metrics like pass@k, which is inadequate for comprehensive repository-level code generation evaluation. In real-world development, code often needs to call predefined modules (functions, classes, or variables) in the repository to align with developer intent. While LLMs may generate code that passes tests, it can result in inefficient implementations or unnecessary duplication of predefined functions, leading to issues like technical debt and code smells (Maldonado and Shihab, 2015; Sierra et al., 2019; Santos et al., 2018).
To address these gaps, we propose RepoExec, a novel benchmark designed to ensure both functional correctness and effective dependency utilization. It offers an executable environment for reliable functional evaluation and incorporates a mechanism to generate high-coverage test cases, enhancing the robustness of functional correctness assessment. Additionally, we also propose Dependency Invocation Rate (DIR), a metric that evaluates the extent to which generated code leverages predefined dependencies, offering a more comprehensive measure of model performance beyond merely passing test cases.
Our experiments with RepoExec offer several key insights into repository-level code generation. Firstly, given varying context levels (full, medium, small), which correspond to the information provided for dependencies—such as full implementation, docstrings, and function signatures—we observe that models achieve the best performance when given full context dependencies while also showing potential with smaller contexts. Secondly, pretrained LLMs and instruction-tuned LLMs exhibit different strengths: pretrained LLMs generally excel in pass@k metrics, while instruction-tuned LLMs perform better on the Dependency Invocation Rate (DIR). This suggests that while pretrained LLMs can generate runnable code, they often struggle with properly utilizing dependencies, which means the code may pass tests but might not be correctly implemented in a repository-level context. On the other hand, instruction-tuned LLMs make better use of contextual information. These findings highlight the importance of using both pass@1 and DIR for a more holistic evaluation, emphasizing the need to improve both functional correctness and context utilization in CodeLLMs.
Furthermore, we investigate two methods to enhance repository-level code generation: multi-round debugging and instruction tuning with dependency contexts. Multi-round debugging with test execution, particularly with models like GPT-3.5 and WizardCoder, significantly boosts both pass@1 and DIR after three rounds. Additionally, fine-tuning on our dependency-enhanced dataset improves these metrics while reducing computational costs. These findings highlight the benefits of executable code testing, instruction tuning, and iterative debugging for improving CodeLLM performance and managing code dependencies more effectively. Besides, it also demonstrates that better context utilization leads to higher pass@k scores (detailed analysis provided in Sections 6.2, 6.3 and Appendix G). In summary, our contributions are as follows:
@@
    We introduce a novel evaluation paradigm for repository-level code generation, assessing both functional correctness and quality factors such as maintainability and adherence to clean code principles through efficient dependency use.
@@
    We introduce RepoExec, a novel benchmark that aligns with our evaluation paradigm, addressing the gaps in existing benchmarks and offering a comprehensive assessment of code generation quality. Additionally, an effective pipeline is introduced within RepoExec from dependency extraction, dynamically generate high-coverage test cases and automatic evaluation with execution and code dependencies. Our pipeline offers practical usage and scalability for the community.
@@
    We release a tool named pydepcall to extract dependencies for all functions within any repository, providing a practical use for advancing research in this domain.
@@
    Our experiments reveal key insights into CodeLLMs’ performance in repository-level code generation. While foundation models show high initial accuracy, instruction-tuned models excel in dependency management. Additionally, multi-round debugging tests further improve performance, enhancing dependency management. Notably, a strong correlation between pass@1 and DIR underscores that better context utilization leads to more functionally correct code.
@@ 2 Related works
Coding-related tasks have been crucial for assessing the performance of Large Language Models (LLMs), with code generation emerging as a primary focus (Chen et al., 2021; Li et al., 2023; Jiang et al., 2024; Touvron et al., 2023; Roziere et al., 2023; Xu et al., 2022; Allal et al., 2023; Nijkamp et al., 2022; Phan et al., 2024a; To et al., 2023). Early benchmarks have been introduced to address this issue Yin et al. (2018); Iyer et al. (2018); Nguyen et al. (2023); Chen et al. (2021); Austin et al. (2021); Hendrycks et al. (2021); however, they often had limited scope or employed weak evaluation approaches. Some benchmarks Yin et al. (2018); Iyer et al. (2018); Nguyen et al. (2023) exhibit domain diversity akin to real-world applications; however, their evaluation methodologies are constrained to match-based metrics, thereby decreasing the reliability of these benchmarks Chen et al. (2021). Meanwhile, benchmarks with reliable evaluation approaches such as HumanEval Chen et al. (2021), MBPP Austin et al. (2021) and APPS Hendrycks et al. (2021) often entail limitations to specific domains like competitive programming. Recently, there have been efforts to extend the domains of generation tasks in various benchmarks. For instance, ExeDS Huang et al. (2022) focuses on data science code generation, while ODEX Wang et al. (2022) serves as an open-domain benchmark for code generation. Besides, all the benchmarks mentioned primarily focus on standalone function generation, lacking consideration for cross-contextual and dependency invocation scenarios.
@@ 3 Evaluation Paradigm
In this section, we outline our paradigm to achieve a more robust and comprehensive evaluation of repository-level code generation. Our paradigm encompasses two key attributes: Functional correctness and Dependency utilization.
Functional correctness:
This evaluation criterion ensures that the code accurately performs its intended tasks and requirements. Specifically, it typically involves using test cases to compare the execution output of the generated code against the expected output for a given input. This criterion has been widely employed for evaluating code generation in numerous studies (Chen et al., 2021; Austin et al., 2021; Zhuo et al., 2024; Li et al., 2024). We follow the well-known automatic metric Pass@k Chen et al. (2021) to measure the functional correctness of generation outputs.
Dependency utilization:
Functional correctness alone cannot fully capture code quality, as it may overlook poor implementations or redundancy. Match-based metrics like BLEU and edit similarity evaluate alignment between generated and reference code at the token level. However, not all tokens impact code quality equally, and alternative implementations can maintain quality despite low similarity scores. Tokens representing called dependencies—such as packages, functions, variables, and classes within the repository—imply effective use of human intent for efficient implementation. Ignoring these dependencies may suggest workaround implementations misaligned with human intent, leading to increased verification and maintenance costs. To assess the models’ ability to utilize provided dependencies in accordance with human intent, we introduce the Dependency Invocation Rate (DIR). This metric represents the percentage of invoked dependencies out of the total number of dependencies provided. Let $D_{g}$ denote the set of identifiers in the generated output, and $D_{s}$ denote the set of provided dependencies extracted from the solution. The Dependency Invocation Rate (DIR) is calculated as follows:
@@
A higher DIR indicates that the model successfully incorporates a larger proportion of the provided dependencies into the generated code, demonstrating a better understanding of the dependencies’ relevance and their intended usage. Conversely, a lower DIR suggests that the model may struggle to identify and utilize the appropriate dependencies, potentially generating code that is less aligned with the human-intent solution. In summary, achieving a high-quality solution requires the generated code to excel in both functional correctness and dependency utilization. Otherwise, it indicates a lack of ability to generate correct solutions or suggests poor implementation practices, including technical debt or code smell issues.
@@ 4 Data Collection Pipeline
Developing an executable benchmark within repository contexts is challenging due to complex setup requirements and frequent lack of clear installation guidelines in repositories. Previous studies (Ding et al., 2024; Shrivastava et al., 2023; Ding et al., 2023; Liao et al., 2023) have often relied on match-based metrics for evaluation, which may not fully capture code functionality. In addition, test cases are essential for assessing code functionality. However, extracting test cases from repositories (Zhang et al., 2023b; Li et al., 2024; Zhang et al., 2024) often relies on available functions and heuristic rules, limiting adaptability and excluding data without existing tests. For instance, Li et al. 2024 found that over 99% of functions were discarded due to the absence of suitable tests. We propose a dataset collection pipeline to ensure that repositories can build executable environments and that test cases are automatically generated. The complete data pipeline is illustrated in Figure 1, with details of the data sources provided in Appendix A.
Table 1: Dataset attributes with different levels of contexts. Cross-file refers to the percentage of problems that involve cross-file dependencies. Tree-sitter is used for tokenization. AVG LC: Average Line Coverage.
@@ 4.1 Functions and Dependencies Extraction
We employ static analysis on call graph to identify dependencies, excluding identifiers that are function parameters or typing objects. Each dependency name is then mapped to its definition within the repository using a repository graph and static analysis tools. We release our tool pydepcall for community usage and provide a brief description in Appendix I.
@@ 4.2 Test case generation
To overcome the limitation of requiring available tests in the repository for evaluation as in previous works, we leverage large language models (LLMs) to generate test cases automatically. Our proposed approach ensures the correctness of created test cases while also controlling and enhancing test coverage. To execute and validate the generated test cases, it is necessary to configure each repository to create an executable environment. Follow Lemieux et al. 2023, we use pipreqs [^1] https://github.com/bndr/pipreqs to identify each project’s dependency packages.
In our test generation process, we conduct two phases corresponding to Correctness Control and Coverage Enhancement. After the test generation process, we exclude samples with a line coverage lower than 40%, as they are insufficient to accurately assess the correctness of the generated code.
@@ 4.2.1 Correctness Control
In this stage, we present the procedure for generating initial test cases and ensure the tests’ correctness. Specifically, we use CodeLlama-13b Roziere et al. (2023) and provide the model with the prompt detailed in Appendix D. The first 20 assertions form the test cases, which are then filtered via syntax and execution checks to ensure correctness.
@@
We use pytest [^2] https://github.com/pytest-dev/pytest to run the generated tests, discarding those with errors except for AssertionError. If an assertion error occurs, the Assertion Fixer resolves it by executing the test with the target function, extracting the output, and updating the assertion. To handle complex return types (e.g., custom objects), we employ pickle [^3] https://docs.python.org/3/library/pickle.html to store results. Additionally, to address flaky tests that are inconsistent due to randomness, each test is executed 10 times for comparison.